# Measuring the Resolved Star Formation Main Sequence in TNG100: Fitting Technique Matters

### v1

@@ 4 Summary & Discussion
We would like to acknowledge the work and documentation provided by the IllustrisTNG team that has made this paper possible. The IllustrisTNG simulations were undertaken with compute time awarded by the Gauss Centre for Supercomputing (GCS) under GCS Large-Scale Projects GCS-ILLU and GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS), as well as on the machines of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany. The additional computational work done for this paper was performed on the Shared Computing Cluster which is administered by Boston University’s Research Computing Services. This work was partially supported by NSF grant AST-2009397. BM acknowledges support by Northeastern University’s Future Faculty Postdoctoral Fellowship Program.

### v3

@@ 4 Summary & Discussion
We thank the anonymous reviewer for their incredibly helpful feedback. We would like to acknowledge the work and documentation provided by the IllustrisTNG team that has made this paper possible. The IllustrisTNG simulations were undertaken with compute time awarded by the Gauss Centre for Supercomputing (GCS) under GCS Large-Scale Projects GCS-ILLU and GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS), as well as on the machines of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany. The additional computational work done for this paper was performed on the Shared Computing Cluster (SCC) which is administered by Boston University’s Research Computing Services. The SCC is located in the Massachusetts Green High Performance Computing Center. This work was partially supported by NSF grant AST-2009397. B.M. acknowledges support by Northeastern University’s Future Faculty Postdoctoral Fellowship Program.